# Soft Theorems in Chern-Simons Matter Theories

### v2

@@ Abstract
I will describe tree level soft theorems for Chern-Simons QED and QCD in 4+1 dimensions. Soft theorems have been studied in various context in the past. Recently, there has been renewed interest in these theorems due to their connection with asymptotic symmetries and gauge invariance. Chern-Simons theories are gauge invariant only up to a boundary term. Hence, they make for an interesting case to study the relationship between gauge invariance and soft theorems. The aim of this work is to include corrections due to the addition of Chern-Simons terms in the gauge field action. I will start with an overview of Chern-Simons theories in 4+1 dimensions. Then I will move on to a discussion on soft theorems in gauge invariant QFT at the subleading order using Sen’s covariantisation procedure. Finally, I will calculate the correction due to the addition of a Chern-Simons term for both QED and QCD.
1 Introduction
Soft theorems are statements about scattering amplitudes of processes involving external states with low energy (soft) gauge bosons. The leading soft theorems relate the amplitudes with soft bosons to those without them by a “universal” soft factor which has a pole divergence in the soft momenta. It has been shown in various previous works that the leading order soft theorems are related to the gauge invariance of the scattering amplitudes via appropriate Ward identities He:2014cra ; Kapec:2015ena ; Kapec:2014zla ; Lysov:2014csa ; Laddha:2017vfh ; Campiglia:2016hvg ; Campiglia:2015qka ; Conde:2016csj ; Hirai:2018ijc ; Mao:2017tey (also a nice review is Strominger:2017zoo ). This naturally leads to the following question: How do soft theorems work for Chern-Simons theories, where the gauge invariance only holds up to a boundary term (for instance, refer to this review Dunne:1998qy ). In this note, we will calculate how the leading and subleading soft theorems get corrected when we introduce a Chern-Simons term in QED and QCD.
Chern-Simons terms are only defined in odd space-time dimensions. In this paper, we will work in 4+1 space-time dimensions instead of the usual 2+1 dimensions since we want to treat the Chern-Simons term perturbatively about the Yang Mills theory. [^1] In D=2+1, the Chern-Simons term goes like $\sim\mathbb{A}d\mathbb{A}+\mathbb{A}^{3}$ which dominates over the Yang Mills action in the infrared. This not the case in higher dimensions where the Chern-Simons term becomes less relevant in the infrared. To state the problem precisely, consider a Yang Mills gauge theory coupled to matter described by the gauge field action $\mathcal{S}_{gauge}$ in 4+1 dimensions. We add a term $\mathcal{S}_{CS}$ such that,
$S=S_{gauge}+S_{CS},$
$S_{gauge}=\int d^{4+1}x\left\{-\frac{1}{4g^{2}}F^{2}-\frac{1}{2g^{2}}\left(\partial.A\right)^{2}+\mathcal{L}_{matter}+\mathcal{L}_{int}\right\},$
$S_{CS}=\kappa\int\mbox{Tr}\left[\mathbb{A}\left(d\mathbb{A}\right)^{2}+c_{4}\mathbb{A}^{3}d\mathbb{A}+c_{5}\mathbb{A}^{5}\right],$
where we have introduced the gauge one form $\mathbb{A}=A_{\mu}dx^{\mu}$ and the product is to be understood as the wedge product $\wedge$. When $\left\{c_{4},c_{5}\right\}=\left\{\frac{3}{2},\frac{3}{5}\right\}$, $\mathcal{S}_{CS}$ becomes a Chern-Simons term Miskovic:2005di and is gauge invariant up to a boundary term [^2] For non-abelian Chern-Simons term, the action is also shifted by a constant which is referred to as the winding number Dunne:1998qy . This does not affect the path integral since for appropriate values of $\kappa$, the winding number is $2\pi$ times an integer and hence leaves $e^{i\mathcal{S}}$ invariant. . Note that if $A_{\mu}$ were abelian, only the first term in $\mathcal{S}_{CS}$ would be non zero and it would be gauge invariant by itself.
In the absence of the Chern-Simons term $S_{CS}$, the subleading soft theorem (for instance, with a soft photon with momentum $k_{\mu}$ and polarisation $e_{\mu}$) takes the form,
where $\varepsilon_{i},p_{i}$ refers to the polarisation and energy of the external finite energy (“hard”) particles and $q_{i}$ refers to the charge of the matter coupling to the photon field. Similar expressions are there for gluons and gravitons as well. The leading term has a $\nicefrac{{1}}{{k}}$ divergence while the subleading term is $\mathcal{O}\left(k^{0}\right)$. For soft photons and gravitons, the leading soft theorems in (1) are shown to be valid to all loop orders. The subleading term gets corrected in $D=2+1$ due to infrared effects, but in $D=4+1$ the subleading soft photon theorem is also valid to all loop orders [^3] I would like to thank Biswajit Sahoo for pointing this out.. For soft gluons, there are infrared corrections at loop level even at the leading order in $D=2+1$. In the case of multiple soft bosons, each boson brings a soft factor of divergence $\nicefrac{{1}}{{k}}$ at leading order. The aim of this note is to calculate how (1) gets corrected at tree level when $\mathcal{S}_{CS}$ is added.
The paper is organized as follows - In section 3, we will review the soft photon and gluon theorems and derive them by covariantising the hard particle action by minimally coupling the soft boson field modes. AtulBhatkar:2018kfi ; Sen:2017nim ; Chakrabarti:2017ltl ; Sen:2017xjn ; Laddha:2017ygw demonstrate it for photons and gravitons. We will generalize it to a generic non-abelian Yang Mills theory with SU(N) gauge invariance and derive the soft gluon theorem. Soft gluon theorems have been previously studied in various ways in Mao:2017tey ; Broedel:2014fsa ; Broedel:2014bza ; Klose:2015xoa . Then in section 4, we will calculate the Feynman rules for the Chern-Simons vertices. In section 5, 6, and 7, we will incorporate them to obtain the Chern-Simons correction to the soft theorems. We will see that the leading soft theorems remain unchanged and the Chern-Simons terms only correct the sub-leading soft theorems.
2 Notations and Conventions
@@
Throughout this paper, I have worked in natural units with $\hbar=c=1$. I have also used the “mostly minus” signature $\left(+---\right)$ for the space-time metric. $\epsilon^{\mu\nu\rho\sigma}$ refers to the Levi-Civita symbol normalised as $\epsilon^{0123}=1$. $\mu$, $\nu$ denote the Lorentz indices while $\alpha$, $\beta$ denote the indices of the representation of the gauge group. The latter will often be omitted to tidy up the equations, in which case the order of multiplication is important. Lorentz indices are raised and lowered with the Minkowski metric $\eta_{\mu\nu}$. Whenever mentioned, the $\gamma^{\mu}$ denote the Dirac matrices satisfying the Clifford algebra $\left\{\gamma^{\mu},\gamma^{\nu}\right\}=2\eta^{\mu\nu}$. We use the Feynman slash notation where $\cancel{b}=b^{\mu}\gamma_{\mu}$ for any four vector $b^{\mu}$. $A_{\mu}$, $a_{\mu}$ are spin-1 gauge-fields in Lorentz gauge with the latter (with small $a$) denoting a soft gauge mode. $F_{\mu\nu}$ is the field strength tensor. We will be working with $SU\left(N\right)$ non-gauge theories for which $a_{\mu}$ etc. are matrices in the representation. They can be expanded in terms of the generators of the gauge group $a_{\mu}\left(x\right)=a_{a\mu}\left(x\right)T_{a}$. The following types of shorthand for index contractions will be present,
@@
3 Subleading Single Soft Theorem from Gauge Invariance
@@
In this section we will summarize the subleading single soft theorem for photons, as given in AtulBhatkar:2018kfi using Sen’s covariantisation procedure Sen:2017nim ; AtulBhatkar:2018kfi ; Chakrabarti:2017ltl ; Sen:2017xjn ; Laddha:2017ygw and generalize it for gluons. The familiar soft theorems (1) get subleading infrared corrections in case of photons and leading infrared corrections in case of gluons. Therefore, we will restrict ourselves to only tree level amplitudes even in this general covariantisation procedure. We start by considering a scattering process for hard particles, with amplitude $G_{n}\left(p_{i}\right)$. We assume that the hard particles in the process are described by a single set of fields $\phi_{\alpha}$, belonging to some reducible representation of the Lorentz group. Let $\mathcal{S}_{hard}\left[\phi\right]$ denote the corresponding action for the hard particle fields. We assume that this action is invariant under the global transformation,
@@
where I’ve used (5) in the second last line of (16) and (8) in the last line of (16). If the soft boson line in $\Gamma^{\left(3\right)}$ is an internal line, then (16) needs to be modified by stripping of the polarisation vector and contracting the index with the appropriate soft propagator. Using (16) and the relations (8), (10), (12) and (13), we can write-
@@
Special Case: Soft Photon Theorem in QED
@@
Special Case: Soft Gluon Theorem in QCD
@@
This is the single soft gluon theorem in QCD.
@@ 4.1 Abelian Case - Chern-Simons QED
Consider again the abelian gauge field $A_{\mu}$. The Chern-Simons term int this case is given by,
$\mathcal{L}$ $=\kappa\int\mathbb{A}\left(d\mathbb{A}\right)^{2}=\kappa\int d^{4+1}x\,\epsilon^{\mu\nu\rho\sigma\delta}A_{\mu}\partial_{\nu}A_{\rho}\partial_{\sigma}A_{\delta},$ (27)
where $\mathcal{L}_{matter}$ is the part of the Lagrangian that contains the matter field. From the last term, we can calculate the Chern-Simons 3-point vertex. The vertex factor in the Feynman rules is again found by symmetrising the Chern-Simons term in external momenta and factoring out the external polarisation factors $\varepsilon_{\mu}\left(k\right)$.
@@ 6 Chern-Simons Correction to Single Soft Gluon Theorem in QCD
where $\widetilde{\Sigma}$ only sums over external gluon legs.
@@ 8 Conclusion
We have shown in (35) and (39) that the Chern-Simons corrections to the soft photon and soft gluon theorems are only subleading in powers of $k$. The leading soft theorems remain unchanged even in the presence of a Chern-Simons term. This is interesting since we can take arbitrary values of $\left\{c_{4},c_{5}\right\}$ which would leave $\mathcal{S}_{CS}$ completely non gauge-invariant. Therefore, we have an example of a non gauge-invariant action that still satisfies the leading soft theorem at the tree level. This means that even though gauge invariance implies the leading soft theorems, the opposite is not true. There are examples of non-gauge invariant theories which still obey the leading soft theorem. I have provided one such example.
@@ Acknowledgments
I would like to thank my PhD advisor, R. Loganayagam, for helping me with the initial idea and in setting up the problem as well as for countless insights. I would also like to thank Biswajit Sahoo and Ashoke Sen for their invaluable feedback on my draft. Finally, I would like to thank my colleagues Godwin Martin, Kaustubh Singhi and Shridhar Vinayak for contributing to various discussions throughout this work. I also acknowledge the support of the Department of Atomic Energy, Government of India, under project no. RTI4001.

### v3

@@ Abstract
We describe tree level soft theorems for Chern-Simons QED and QCD in 4+1 dimensions. Soft theorems have been studied in various context in the past. The universal behavior of the leading and subleading soft theorems is understood as a consequence of the local on-shell invariance of the scattering amplitudes. Chern-Simons terms are gauge invariant only up to a boundary term. Hence, they make for an interesting case to study the relationship between gauge invariance and soft theorems. In this paper, we show that the subleading soft theorems in QED and QCD get corrected when Chern-Simons terms are introduced and also derive those corrections. We start with a discussion on soft theorems as a consequence of gauge invariance. Then we move on to derive the Chern-Simons vertices. Finally, we calculate the correction to soft theorems due to the addition of Chern-Simons terms for both QED and QCD.
Notations and Conventions
@@
Throughout this paper, I have worked in natural units with $\hbar=c=1$. I have also used the “mostly minus” signature $\left(+---\right)$ for the space-time metric. $\epsilon^{\mu\nu\rho\sigma}$ refers to the Levi-Civita symbol normalized as $\epsilon^{0123}=1$. $\mu$, $\nu$ denote the Lorentz indices while $\alpha$, $\beta$ denote the indices of the representation of the gauge group. The latter will often be omitted to tidy up the equations, in which case the order of multiplication is important. Lorentz indices are raised and lowered with the Minkowski metric $\eta_{\mu\nu}$. Whenever mentioned, the $\gamma^{\mu}$ denote the Dirac matrices satisfying the Clifford algebra $\left\{\gamma^{\mu},\gamma^{\nu}\right\}=2\eta^{\mu\nu}$. We use the Feynman slash notation where $\cancel{b}=b^{\mu}\gamma_{\mu}$ for any four vector $b^{\mu}$. $A_{\mu}$, $a_{\mu}$ are spin-1 gauge-fields in Lorentz gauge with the latter (with small $a$) denoting a soft gauge mode. $F_{\mu\nu}$ is the field strength tensor. We will be working with SU($N$) gauge theories for which $a_{\mu}$ etc. are matrices in the representation. They can be expanded in terms of the generators of the gauge group $a_{\mu}\left(x\right)=a_{a\mu}\left(x\right)T_{a}$. The following types of shorthand for index contractions will be present,
@@
1 Introduction
Soft theorems are statements about scattering amplitudes of processes involving external states with low energy (soft) gauge bosons. They relate the amplitudes with external soft gauge bosons to those without them via a soft operator. This soft operator is given as an expansion in the soft momenta of the external gauge bosons. The leading term in the expansion has a pole divergence in the soft momenta Gell-Mann:1954wra; Low:1954kd; Low:1958sn; Burnett:1967km; Weinberg:1964ew; Weinberg:1965nx. The leading and subleading terms in the soft operator have a universal form which is understood to be a consequence of local on-shell gauge invariance of the scattering amplitudes Low:1954kd; Low:1958sn; Weinberg:1964ew; Burnett:1967km; Bern:2014vva. This naturally leads to the following question: does this statement apply to gauge theories with Chern-Simons terms as well? Unlike the usual gauge invariant terms in the QED or QCD action, the Chern-Simons terms are gauge invariant only up to a boundary term Chern:1974ft; Dunne:1998qy. Therefore, they do not modify the Ward identity corresponding to small gauge transformations. Naturally then it is interesting to ask whether they modify the universal form of the leading or subleading soft theorems. In this paper, we answer this question. We calculate how the soft theorems get corrected at the tree level when we introduce a Chern-Simons term in QED and QCD.
Chern-Simons terms are only defined in odd space-time dimensions. They are usually studied in 2+1 space-time dimensions. We are instead working in 4+1 dimensions because we want to treat the Chern-Simons term perturbatively about the usual gauge theory of QED or QCD. [^1] In D=2+1, the Chern-Simons term goes like $\sim\mathbb{A}d\mathbb{A}+\mathbb{A}^{3}$ which dominates over the Yang Mills action in the infrared. This not the case in higher dimensions where the Chern-Simons term becomes less relevant in the infrared. Hence, we consider a gauge field coupled to matter in $D=4+1$,
$S_{gauge}=\int d^{4+1}x\left\{-\frac{1}{4g^{2}}F^{2}-\frac{1}{2g^{2}}\left(\partial.A\right)^{2}+\mathcal{L}_{matter}+\mathcal{L}_{int}\right\},$ (1)
where $A_{\mu}$ is the gauge field and $F_{\mu\nu}$ is the appropriate field strength tensor. In the case of QED, the gauge field is abelian and hence we add the abelian Chern-Simons term given by,
$\mathcal{S}_{CS}$ $=\kappa\int\mathbb{A}\left(d\mathbb{A}\right)^{2}=\kappa\int d^{4+1}x\,\epsilon^{\mu\nu\rho\sigma\delta}A_{\mu}\partial_{\nu}A_{\rho}\partial_{\sigma}A_{\delta}.$ (2)
where we have introduced the gauge one form $\mathbb{A}=A_{\mu}dx^{\mu}$ and the product is to be understood as the wedge product $\wedge$. For QCD, we need to add the non-abelian Chern-Simons form in 4+1 dimensions,
$S_{CS}=\kappa\int\mbox{Tr}\left[\mathbb{A}\left(d\mathbb{A}\right)^{2}+c_{4}\mathbb{A}^{3}d\mathbb{A}+c_{5}\mathbb{A}^{5}\right],$ (3)
When $\left\{c_{4},c_{5}\right\}=\left\{\frac{3}{2},\frac{3}{5}\right\}$, $\mathcal{S}_{CS}$ becomes a Chern-Simons term Miskovic:2005di and is gauge invariant up to a boundary term [^2] For non-abelian Chern-Simons term, under large gauge transformations the action is also shifted by a constant which is referred to as the winding number Dunne:1998qy. This does not affect the path integral since for appropriate discrete values of $\kappa$, the winding number is $2\pi$ times an integer and hence leaves $e^{i\mathcal{S}}$ invariant. . The total action is then given by the sum of the two,
$S=S_{gauge}+S_{CS},$
In the absence of the Chern-Simons term $S_{CS}$, the subleading soft theorem (for instance, for an external soft photon with momentum $k_{\mu}$ and polarization $e_{\mu}$) takes the form,
where $\varepsilon_{i},p_{i}$ refers to the polarization and energy of the external finite energy (“hard”) particles and $q_{i}$ refers to the charge of the matter coupling to the photon field. Similar expressions are there for gluons and gravitons as well. The leading term has a $\nicefrac{{1}}{{k}}$ pole divergence while the subleading term is $\mathcal{O}\left(k^{0}\right)$. The soft photon theorem at leading order in $k$ does not get infrared corrections at loop level Bern:2014vva; Bern:2014oka; He:2014bga; He:2014cra. In $D=2+1$, the subleading term gets corrected due to infrared effects at loop level, but in $D=4+1$ the subleading soft photon theorem is also valid to all loop orders. The soft gluon theorem in $D=2+1$ admits infrared corrections at loop level even at the leading order but in $D=4+1$, the theory becomes IR finite and corrections occur only at subleading order. In the case of multiple soft gauge bosons, each gauge boson brings a soft factor of divergence $\nicefrac{{1}}{{k}}$ at leading order.
As mentioned before, the leading and subleading behavior in (4) is universal and can be understood as a consequence of local on-shell gauge invariance of the scattering amplitudes. This was shown for photons in Low:1958sn and for gluons in Bern:2014vva; Broedel:2014fsa. For gravitons, even the second subleading behaviour (not stated here) is universal and can be obtained from gauge invariance of the amplitudes Weinberg:1964ew; Weinberg:1965nx; Bern:2014vva; Broedel:2014fsa. Sen’s covaraintisation procedure Sen:2017nim; Sen:2017xjn; AtulBhatkar:2018kfi; Laddha:2017ygw; Chakrabarti:2017ltl utilizes gauge invariance to derive this universal behavior by giving a covariantisation prescription to calculate the 3-point vertex (see $\Gamma^{\left(3\right)}\left(e,k;p,-p-k\right)$ in figure 1) for the leading order diagram. Here, one writes an action involving only hard particle fields and covariantises it by minimally coupling the gauge field mode for the soft gauge boson. This is described in detail in section 2. Such a covariantisation prescription does not apply to Chern-Simons terms because even though the small gauge Ward identity is still valid, the Chern-Simons terms are not minimally coupled to the matter. Hence, their contribution needs to be accounted for separately in the 3-point vertex, $\Gamma^{\left(3\right)}\left(e,k;p,-p-k\right)$.
Furthermore, we show that gauge invariance is a sufficient but not necessary condition for having the universal form of the leading soft theorems. For this, we will take $c_{4}$ and $c_{5}$ in (3) to be arbitrary (and hence make $\mathcal{S}_{CS}$ non gauge-invariant) and show that the universal form of the leading soft theorem still holds and only the subleading term gets corrected.
Hence, ultimately the aim of this paper is to calculate how (4) gets corrected at tree level when a Chern-Simons term $\mathcal{S}_{CS}$ is added to the gauge theory.
The paper is organized as follows - in section §2, we review the soft photon and gluon theorems and derive them by covariantising the hard particle action by minimally coupling the soft gauge field modes. AtulBhatkar:2018kfi; Sen:2017nim; Chakrabarti:2017ltl; Sen:2017xjn; Laddha:2017ygw demonstrate it for soft photons and gravitons. We generalize it to a generic non-abelian Yang Mills theory with SU($N$) gauge invariance and derive the soft gluon theorem. Soft gluon theorems have been previously studied in various ways in Bern:2014vva; Mao:2017tey; Broedel:2014fsa; Broedel:2014bza; Klose:2015xoa. Then in section §3, we calculate the Feynman rules for the Chern-Simons vertices in $D=4+1$. In section §4, §5 and §6 we incorporate them to obtain the Chern-Simons correction to the soft theorems. We see that the leading soft theorems remain unchanged and the Chern-Simons terms only correct the subleading soft theorems.
2 Subleading Soft Theorem from Sen’s Covariantisation Procedure
@@
In this section we will summarize the subleading single soft theorem for photons, as given in AtulBhatkar:2018kfi using Sen’s covariantisation procedure Sen:2017nim; AtulBhatkar:2018kfi; Chakrabarti:2017ltl; Sen:2017xjn; Laddha:2017ygw and generalize it for gluons. For gluons, the amplitudes and the corresponding diagrams will be implicitly assumed to be color ordered. At loop level, the familiar soft theorems (4) get subleading infrared corrections in case of photons and leading infrared corrections in case of gluons. Therefore, we will restrict ourselves to only tree level amplitudes even in this general covariantisation procedure. We start by considering a scattering process for hard particles, with amplitude $G_{n}\left(p_{i}\right)$. We assume that the hard particles in the process are described by a single set of fields $\phi_{\alpha}$, belonging to some reducible representation of the Lorentz group. Let $\mathcal{S}_{hard}\left[\phi\right]$ denote the corresponding action for the hard particle fields. We assume that this action is invariant under the global transformation,
@@
where I’ve used (8) in the second last line of (19) and (11) in the last line of (19). If the soft boson line in $\Gamma^{\left(3\right)}$ is an internal line, then (19) needs to be modified by stripping of the polarization vector and contracting the index with the appropriate soft propagator. Using (19) and the relations (11), (13), (15) and (16), we can write-
@@
Soft Photon Theorem in QED
@@
Soft Gluon Theorem in QCD
@@
This is the single soft gluon theorem in QCD. Note that the sum will only contain terms that are in accordance with the appropriate color ordering.
@@ 3.1 Abelian Case - Chern-Simons QED
Consider the abelian gauge field $A_{\mu}$. The Chern-Simons term in this case is given by,
$\mathcal{S}_{CS}$ $=\kappa\int\mathbb{A}\left(d\mathbb{A}\right)^{2}=\kappa\int d^{4+1}x\,\epsilon^{\mu\nu\rho\sigma\delta}A_{\mu}\partial_{\nu}A_{\rho}\partial_{\sigma}A_{\delta}.$ (30)
We can calculate the Chern-Simons 3-point vertex. The vertex factor in the Feynman rules is again found by symmetrizing the Chern-Simons term in external momenta and factoring out the external polarization factors $\varepsilon_{\mu}\left(k\right)$.
@@ 5 Chern-Simons Correction to Single Soft Gluon Theorem in QCD
where $\widetilde{\Sigma}$ only sums over external gluon legs. Again, we assume that the sums only contain terms that are in accordance with the color ordering of the amplitude.
@@ 7 Conclusion
We have shown in (38) and (42) that the Chern-Simons corrections to the soft photon and soft gluon theorems are subleading in powers of soft momenta $k$. The leading soft theorems remain unchanged even in the presence of a Chern-Simons term. This is interesting for two reasons -
(1) the subleading theorem gets modified when Chern-Simons terms are added and therefore, it is no longer universal even though the theory is still gauge-invariant. Also, such terms cannot be obtained from Sen’s covariantisation procedure and need to be added separately.
(2) we can in principle take arbitrary values of $\left\{c_{4},c_{5}\right\}$ which would leave $\mathcal{S}_{CS}$ completely non gauge-invariant and still the amplitudes will obey the "universal" leading soft theorem. Therefore, we have an example of a non gauge-invariant action that still satisfies the usual universal leading soft theorem at the tree level. This means that even though gauge invariance implies the universal form of leading soft theorems, the opposite is not true. There are examples of non-gauge invariant theories which still obey the same leading soft theorem. I have provided one such example.
@@ Acknowledgments
I would like to thank my PhD advisor, R. Loganayagam, for helping me with the initial idea and in setting up the problem as well as for countless valuable insights. I would also like to thank Biswajit Sahoo and Ashoke Sen for their invaluable feedback on my draft. Finally, I would like to thank my colleagues Godwin Martin, Kaustubh Singhi and Shridhar Vinayak for contributing to various discussions throughout this work. I also acknowledge the support of the Department of Atomic Energy, Government of India, under project no. RTI4001.